OpenStudy (anonymous):

The probability of a computer chip failing is 2% . if 100 computers are sold, what is the probability of 2 computer chips failing?

OpenStudy (caozeyuan):

(2%)^2*(98%)^98

OpenStudy (anonymous):

what is it as a percent ? @caozeyuan

OpenStudy (caozeyuan):

you can try calculating on your caculator, I'll do that in a sec

OpenStudy (anonymous):

i did, but it doesn match up to any of the answers.

OpenStudy (anonymous):

the possible answers are 8.1 % , 2.7% and 27.3% and 12.1 %

OpenStudy (caozeyuan):

shoot!

OpenStudy (caozeyuan):

Do you have the key? if you do, check if 27.3% is right

OpenStudy (anonymous):

i dont have the key.

OpenStudy (kropot72):

The solution is found from the binomial distribution, as follows: \[\large P(exactly\ 2\ fail)=100C2\times(0.02)^{2}\times(0.98)^{98}=you\ can\ calculate\]

OpenStudy (kropot72):

100C2 = 4950. Therefore the required probability is given by: \[\large 4950\times(0.02)^{2}\times(0.98)^{98}=you\ can\ calculate\] Note: Multiply the result by 100 to give the answer as a percentage, as required by the answer choices.

OpenStudy (caozeyuan):

I forgot my 100C2 cuz I am done with stats 2 years ago, and yeasterday I am done with College Math completely

OpenStudy (anonymous):

27.3% ? @kropot72

OpenStudy (kropot72):

Yes, that is correct.

OpenStudy (anonymous):

another question you might be able to help with. ::: two marbles are to be pulled from a bag that contains 2 yellow marbles, 4 green marbles, and 6 orange marbles. After the first marble is drawn, it is not replaced. WHat is the probability that both marbles will be the same color? I know the formula looks something like (6*4*2)/(***)=

OpenStudy (anonymous):

@kropot72 ^^^

OpenStudy (kropot72):

\[\large P(2\ yellow)=\frac{2}{12}\times\frac{1}{11}\] \[\large P(2\ green)=\frac{4}{12}\times\frac{3}{11}\] \[\large P(2\ orange)=\frac{6}{12}\times\frac{5}{11}\] \[\large P(2\ same\ colour)=\frac{1}{132}(2+12+30)=you\ can\ calculate\]

OpenStudy (anonymous):

0.333 ?

OpenStudy (anonymous):

@kropot72 ^^

OpenStudy (kropot72):

Yes, 0.33333. or 1/3.

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

are you good with deviation? @kropot72

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

"are you good with deviation?" Please post as new question.